OpenStudy (anonymous):

Evaluate the limit lim (√(9+4x^(2)))/(11+10x) x->INF

OpenStudy (turingtest):

\[\lim_{x\to\infty}{\sqrt{9+4x^2}\over11+10x^2}\]correct?

OpenStudy (anonymous):

(11+10x)

OpenStudy (turingtest):

\[\lim_{x\to\infty}{\sqrt{9+4x^2}\over11+10x}\]

OpenStudy (anonymous):

yup

OpenStudy (turingtest):

much easier... divide top and bottom by x, then take the limit

OpenStudy (anonymous):

so its √(9+4x^)/(11+10)?

OpenStudy (anonymous):

√(9+4x)/(11+10)?

OpenStudy (turingtest):

you must be more careful with the numerator

OpenStudy (turingtest):

what is\[{\sqrt{9+4x^2}\over x}\]simplified?

OpenStudy (anonymous):

9+4x?

OpenStudy (turingtest):

1) where did the square root go? 2) how does the square root sign affect trying to divide by x?

OpenStudy (turingtest):

the answers: 1) the square root must stay 2) in order to divide this by x that means we must square it, to get it into the radical

OpenStudy (anonymous):

square the numerator

OpenStudy (anonymous):

?

OpenStudy (anonymous):

9+4x^2

OpenStudy (turingtest):

you have to square x to put it into the radical, just as it works the other way....

OpenStudy (turingtest):

\[\sqrt{\frac a{x^2}}=\frac1x\sqrt{a}\]so it works the other way too

OpenStudy (turingtest):

\[{\sqrt a\over x}=\sqrt {a\over{x^2}}\]

OpenStudy (turingtest):

so we have\[{\sqrt{9+4x^2}\over x}=\sqrt{{9+4x^2\over x^2}}=\sqrt{\frac9{x^2}+4}\]for the numerator

OpenStudy (anonymous):

why did you square x?

OpenStudy (turingtest):

it's perhaps best seen by example...

OpenStudy (turingtest):

\[{\sqrt {16}\over2}=\sqrt{\frac{16}{2^2}}=\sqrt{\frac{16}4}=2\]\[{\sqrt {25}\over5}=\sqrt{\frac{25}{5^2}}=\sqrt{\frac{16}4}=1\]

OpenStudy (turingtest):

etc.

OpenStudy (turingtest):

try it with uglier numbers if you like, I don't want to type it all out

OpenStudy (turingtest):

in principle it works the same as simplifying\[\sqrt{24}\]if you know how to do that:\[\sqrt{24}=\sqrt4\cdot6=2\sqrt6\]

OpenStudy (turingtest):

only here we are doing it with division

OpenStudy (turingtest):

\[\sqrt{\frac{50}4}=\frac{\sqrt{50}}2=\frac{2\sqrt5}2=\sqrt5 \]

OpenStudy (turingtest):

I'm sorry, but I'm a bit too tired to go into a more detailed explanation. I hope you can meditate on this and decipher it.

OpenStudy (anonymous):

yup i got it thank you

OpenStudy (turingtest):

well what I wrote above was wrong, should have been:\[\sqrt{\frac{50}4}=\frac{\sqrt{50}}2=\frac{5\sqrt2}2\]told you I'm tired. later!

OpenStudy (anonymous):

hhhh thanks

OpenStudy (anonymous):

2/10 got it